Suppose one of the Global Positioning System satellites has a speed of 4.10 km/s at perigee and a speed of 3.20 km/s at apogee.

If the distance from the center of the Earth to the satellite at perigee is 2.66×104 km , what is the corresponding distance at apogee?
Answer #1

La = hp mrava = Msp Up (8) (3.20 69/8) = (2-66 x 100 km) (4.10 km/&) Ira = 3.4 x 104 km